# Otter-Knowledge: benchmarks of multimodal knowledge graph representation learning from different sources for drug discovery

### v1

@@ 4 Experiments
We summarise the main results from our experiments and further details are available in the appendices.
@@
4.1 Results and discussion
4.2 Benchmark tasks and downstream models
Downstream benchmarks
We use the TDC [12] benchmark datasets regarding drug-target binding affinity for evaluation. The DTI-DG dataset has a leaderboard with the state-of-the-art metrics reported for different methods. The data was temporally split based on patent application dates making this dataset suitable for method generalization evaluation. On the other hand, the DAVIS and the KIBA datasets have a random split, together with two more splits based on target or drug. The latter splits help validate the learning methods against new drugs/proteins never seen during training.
Downstream models
The TDC framework adapts the DeepDTA framework [21] to learn drug/protein features using separated convolution layers before concatenating transformed features for binding affinity prediction. We discovered that the given architecture is not effective in the context where both drug and protein embeddings are already given as input. With the original TDC framework, the baseline used in Table 3, based on ESM embeddings for proteins and Morgan fingerprints for molecules SMILES, gives much worse results (0.539) than using an architecture that directly concatenates protein and drug embeddings for further transformation (0.569) on the DTI DG dataset. Therefore, we create a new model architecture for the evaluation of the pretrained embeddings. Besides the network that transforms the ESM and Morgan fingerprint, we create a parallel network that transforms the GNN embeddings. Both networks predict the binding affinity and the final prediction is the sum of two predictions (for details about the network size refer to the Appendices).
Ensemble learning
The pretrained representation is sensitive to various factors, such as the chosen objectives for the Graph Neural Network (GNN) and the specific graphs used for training. Additionally, combining all the datasets into a single large graph requires substantial computational resources and poses challenges in aligning different databases. In this study, we propose a simple approach that involves combining pretrained embeddings obtained from different settings and datasets. This allows us to train multiple GNN models simultaneously without the need to merge all the datasets into a single location, resulting in savings in computational resources and effort required for database alignment. We create a linear ensemble by assigning equal weights to individual downstream models, each trained with separate pretrained embeddings given by individual GNN models.
@@ 4.2 Benchmark tasks and downstream models
DUDe Otter DistMult 0.577 0.805 0.573 0.132 0.857 0.650 0.607
@@
PrimeKG Otter DistMult 0.575 0.806 0.571 0.162 0.858 0.611 0.617
@@
STITCH Otter DistMult 0.575 0.808 0.573 0.138 0.859 0.615 0.603
@@
Datasets (UBC) DTI DG DAVIS KIBA
@@
Table 4: Information flow control and noisy links results for UBC for different scoring functions. The table results should be compared with the results in Table 3 (UBC). The evaluation metrics is Pearson correlation (higher is better). N (noisy links); C (no flow control); R (regression).
@@
In Table 3, we present the outcomes of the Otter-Knowledge models, which were pretrained on graphs generated from Uniprot (U), BindingDB (B), ChemBL (C), DUDe, PrimeKG and STITCH with three different training objectives: TransE, DistMult, and binary classifier respectively. Also, and as described in Section 2.3, we control the information propagated to the Drug/Protein entities, and manually handpick a subset of links from each database that are relevant to drug discovery. In all of these methods, we started with the initial embeddings of sequences using ESM-1b models, and Morgan fingerprints were utilized for SMILES, we call this baseline method vanilla representation as oppose to methods with knowledge enhanced representation. The embeddings were then fine-tuned with knowledge graphs. Our results demonstrate that Otter-Knowledge outperforms the baseline vanilla representation without the enhanced knowledge from the graphs. Notably, a significant improvement was observed when we created an ensemble of 12 models trained on UDC and DUDe, PrimeKG and STITCH. We achieved state-of-the-art (SOTA) results on the leaderboard of the DTI DG dataset. However, for the KIBA dataset with drug split, the improvements were not substantial. As indicated in Table 2, the KIBA dataset consists of only 68 drugs. The limited number of drugs makes this specific split particularly challenging, and we consider it an open challenge for future research.
@@
Table 4 shows the results of Otter-Knowledge for UBC when (i) we do not control the information that is propagated to Drug/Protein entities, (ii) we do not cherry-pick a subset of links from each database that are relevant to the downstream task, (iii) regression for numerical data properties is added to the objective in addition to link prediction. Observe from the table that the results are similar to the results in Table 3, with minor variations across different scoring functions and datasets. Notably, Otter Classifier with noisy links (N) and no information flow control (C) achieves comparable or even better performance than when we cherry-pick links and control the flow of information (Table 3). These minor variations suggest that the embeddings computed by the GNN are resilient to noisy triples that are not directly relevant to the downstream tasks. Adding regression objectives and information flow control does not provide a significant result improvement.
@@
The base of the Otter-Knowledge is to leverage the pre-trained representations, before enhancing them with additional knowledge. Thus, the initial embeddings computed for the SMILES and sequences have an impact in the results. Table 5 shows the results of Otter-Knowledge using a ClassifierHead, trained on UDB for 25 epochs with different drug initial representations. We can see that the MolFormer does not give superior results compared to Morgan-fingerprint, there is room for improvement regarding learned representation over simple fingerprint-based approaches for small molecules.
@@
Table 5: Impact of different modalities for drugs on the UDB datasets
@@ 5 Related work
[…]
@@
LMs do not consider existent extensive knowledge, in the form of manually curated functional and structural annotations, in human-curated domain knowledge bases and effectively leveraging all this available factual knowledge to enhance representation learning is still an open research challenge. Nonetheless, prior research indicates that it can improve results in downstream learning tasks. OntoProtein [38] fine-tuned a PLM by reconstructing masked amino acids while minimizing the embedding distance between the contextual representation of proteins and associated gene ontology functional annotations [3]. For this purpose they built ProteinKG25, a KG consisting of  600k protein sequences and nearly five million triples. Their results show that the representations obtained where useful for classification tasks such as protein-protein interaction type, protein function, and contact prediction; but underperformed in regression tasks like homology, fluorescence, and stability.
KeAP [39], on the other hand, claims to explore a more granular token-level approach, where non-masked amino acids iteratively query the associated knowledge tokens to extract helpful information (from the Gene ontology) for restoring masked amino acids via cross-attention. The training process is guided only by the mask token reconstruction objective, while OntoProtein uses both contrastive learning and masked modelling simultaneously. KeAP, also trained on ProteinKG25 [38] and it outperforms OntoProtein on 9 dowstream tasks, including contact, protein-protein interaction type, homology, stability, and protein-protein binding affinity prediction.
@@
Graphs are a natural way to represent molecular interactions, signalling pathways, and disease co-morbidities. They can also be used for representation learning as they allow for the distillation of high-dimensional information about a node’s neighborhood into low-dimensional vector spaces. The training objective of these representations is that similar neighborhoods are embedded close to each other in the vector space. Optimised representations can then be used to train downstream models to predict properties of specific nodes (e.g., protein function), as well as, novel associations between nodes (e.g., drug-target interactions). An overview on graph representation learning in biomedicine can be found in [14]. State of the art approaches have shown that incorporating multiple knowledge sources improves downstream performance (e.g., DTiGEMS+ [35] formulates the prediction of DTIs as a link prediction problem in an heterogeneous graph, or TxGNN for predicting drug indications/contraindications for rare diseases [10]).
@@ 6 Conclusion and future work
In this paper, we studies the representation learning problems for multimodal knowledge graphs fused from multiple sources. Our study serves as a foundation for future investigations in this area with the release of curated datasets from different sources in the format that is ready for studying pretraining methods. Additionally, we have made available pre-trained models that have been constructed using a comprehensive collection of open data. These models can be utilized to acquire representations specifically tailored for drug discovery applications. Furthermore, they can serve as benchmarks for comparing and evaluating against other baseline representation learning techniques. We have also made the standard evaluation framework for assessing pretrained representations in drug-target binding affinity prediction publicly available. Furthermore, we conducted a thorough analysis of various representation learning methods using three well-established benchmark datasets for drug-target binding affinity prediction. Notably, our approach achieved state-of-the-art results on the TDC DG dataset, demonstrating its superiority over existing methods.
Our study establishes the foundation for exploring the learning of multi-modal knowledge representation. Nevertheless, numerous unresolved research questions persist. Firstly, the incorporation of additional modalities, such as the 3D structure of molecules or proteins, can provide valuable insights for representation learning. Secondly, the challenge lies in effectively handling a vast number of databases, where aligning them is not a straightforward task. Developing a learning approach capable of accommodating the dynamic input schema from diverse sources is a crucial problem to address. Finally, evaluating generasibility of the learned graph representation for further (predictive or generative) downstream tasks and use cases and having a more robust learning methods for generalizing the learned representation to multiple tasks under data distribution shift is an important research topic.
@@ 7 Limitations
Due to license restriction of some datasets we only release the datasets with non-commercial licenses and the pretrained models that were built on these datasets. The datasets do not include 3D structure of proteins/drugs which can be an interesting modalities for future work.
@@ Appendix A Appendix
A.1 Experimental settings for upstream/downstream tasks
The downstream models comprise of two parallel structures. In the first architecture, the initial embeddings of drugs and proteins, obtained using the Morgan-fingerprint and ESM-1b models, are concatenated. These embeddings undergo a transformation through two linear layers and RELU activation before predicting the binding affinity. The second architecture involves the concatenation of GNN embeddings of drugs and proteins. Similarly, these embeddings are transformed using two linear layers and RELU activation, followed by predicting the binding affinity.
@@ A.1 Experimental settings for upstream/downstream tasks
The results presented in Table 3 rely on pretrained models that have restricted the information they can access for Protein/Drug entities. These entities are only allowed to utilize amino acid sequences and smiles. Inference tasks follow the pretraining tasks to acquire pretrained GNN embeddings for drugs/proteins in the downstream tasks. Even so Table 4 shows that information flow controls to drugs/proteins is not necessary to achieve good performance on the downstream tasks.
Learning rate has an impact on the downstream tasks, especially for KIBA and DAVIS. Figure 2 and Figure 3 show the sensitivity of setting of learning rate and its impact on the downstream model performance. We have tried different value of learning rates including 5e-4, 5e-5 and 5e-6. The results reported in the paper are based on fixing the learning rate to 5e-4. This value is set as the default learning rate for both the baseline and proposed methods in the work.
@@ A.2 Model release
Table 7 provides a summary of the models. Each model’s name reflects the dataset it was trained on and the scoring function employed during pretraining. For instance, the model otter_ubc_transe denotes the one trained on the Uniprot, BindingDB, and ChemBL (UBC) dataset using the TransE scoring function. All models underwent training for either 35 epochs or within a maximum time frame of 24 hours.
@@
otter_ubc_transe UBC TransE MIT
otter_ubc_distmult UBC DistMult MIT
otter_ubc_classifier UBC Classifier MIT
otter_dude_transe DUDe TransE MIT
otter_dude_distmult DUDe DistMult MIT
otter_dude_classifier DUDe Classifier MIT
otter_primekg_transe PrimeKG TransE MIT
otter_primekg_distmult PrimeKG DistMult MIT
otter_primekg_classifier PrimeKG Classifier MIT
otter_stitch_transe Stitch TransE MIT
otter_stitch_distmult Stitch DistMult MIT
otter_stitch_classifier Stitch Classifier MIT

### v3

@@ 4 Experiments
Here we summarise the main results from some of our experiments demonstrating the benefits of our multimodal KGs. Further details are available in the appendices.
4.1 Benchmark tasks and downstream models
Downstream benchmarks
To evaluate the performance of the proposed approach on drug-target binding affinity prediction task we use three datasets: DTI DG, DAVIS and KIBA, which are available from the TDC [12] benchmark. The DTI DG dataset features a leaderboard with the state-of-the-art metrics reported for different methods. The dataset’s temporal split, based on patent application dates, makes this dataset suitable for evaluating method generalization. In contrast, the DAVIS and KIBA datasets employ random splits, including two additional splits based on target or drug. These latter splits assess learning methods with new drugs/proteins.
Downstream models
TDC framework adapts the DeepDTA approach [21] for drug protein binding affinity prediction model referred to as the downstream model. It consists of 3 blocks: i) convolution block which performs a separate feature learning of input drugs and proteins, ii) concatenation which concatenates drugs and proteins features, and iii) transformation which executes the final transformation. We observed that the given architecture is not optimal when both drug and protein embeddings are already given as input. For this, we compared the original downstream model against a simplified variant that employs only concatenation and transformation blocks. When employing the original downstream model with ESM embeddings for proteins and Morgan fingerprints for molecules SMILES as inputs, we obtained inferior results (0.538, the baseline used in Table 3) compared to using the simplified downstream model (0.569) on the DTI DG dataset. This motivated us to created a new model architecture for evaluating the pretrained embeddings. In addition to the network that concatenates and transforms the ESM and Morgan fingerprints, we added a parallel network that concatenates and transforms the GNN embeddings. The outputs from both networks are summed up and returned as the final binding affinity prediction. We note that the new downstream model uses the same mean squared error loss as the original downstream model (for details about the network size refer to the Appendices).
Ensemble learning
The pretrained representation is sensitive to various factors, such as the chosen objectives for the GNN and the specific graphs used for training. Additionally, combining all the datasets into a single large graph requires substantial computational resources and poses challenges in aligning different databases. In this study, we propose a simple approach that involves combining pretrained embeddings obtained from different settings and KGs. This allows us to train multiple GNN models simultaneously without the need to merge all KGs into a single location, resulting in savings in computational resources and effort required for data and schema alignment. We create a linear ensemble by assigning equal weights to individual downstream models, each trained with separate pretrained embeddings given by individual GNN models.
@@ 4.1 Benchmark tasks and downstream models
Otter-DUDe Otter DistMult 0.577 0.805 0.573 0.132 0.857 0.650 0.607
@@
Otter-PrimeKG Otter DistMult 0.575 0.806 0.571 0.162 0.858 0.611 0.617
@@
Otter-STITCH Otter DistMult 0.575 0.808 0.573 0.138 0.859 0.615 0.603
@@
Datasets (Otter-UBC) DTI DG DAVIS KIBA
@@
Table 4: Information flow control and noisy links results for Otter-UBC for different scoring functions. The table results should be compared with the results in Table 3 (Otter-UBC). The evaluation metrics is Pearson correlation (higher is better). N (noisy links); C (no flow control); R (regression).
4.2 Results and discussion
@@
In Table 3, we present the outcomes of the Otter-Knowledge models, which were pretrained on graphs generated from Uniprot (U), BindingDB (B), ChemBL (C), DUDe, PrimeKG and STITCH with three different training objectives: TransE, DistMult, and binary classifier respectively. Also, as described in Section 2.3, we control the information propagated to the Drug/Protein entities, and manually handpick a subset of links from each database that are relevant to drug discovery. In all these methods, we started with the initial embeddings of sequences using ESM-1b models, while Morgan fingerprints were utilized for SMILES. We call this baseline method vanilla representation as opposed to methods utilizing knowledge-enhanced representations. The embeddings were then fine-tuned with knowledge graphs. Our results demonstrate that Otter-Knowledge outperforms the baseline vanilla representation that lacks the enhanced knowledge from the graphs. Notably, a significant improvement was observed when we created an ensemble of 12 models trained on Otter-UBC, Otter-DUDe, Otter-PrimeKG and Otter-STITCH. We achieved state-of-the-art results on the leaderboard of the DTI DG dataset (Table 3). However, for the KIBA dataset with drug split, the improvements were not substantial. As indicated in Table 4, the KIBA dataset consists of only 68 drugs. The limited number of drugs makes this specific split particularly challenging, and we consider it an open challenge for future research.
@@
Given that drug discovery predominantly involves novel entities with significantly fewer known attributes that those known during pretraining, it is important to assess if the pretraining approach can exhibit robust generalization despite the discrepancy in available data between pretraining and fine-tuning stages. Table 4 shows the results of Otter-Knowledge for Otter-UBC when (i) we do not control the information that is propagated to Drug/Protein entities ("C" in Table 4), (ii) we do not cherry-pick a subset of links from each database that are relevant to the downstream task ("N" in Table 4), (iii) regression for numerical data properties is added to the objective in addition to link prediction ("R" in Table 4). Observe from the table that the results are similar to the results in Table 3, with minor variations across different scoring functions and datasets. Notably, Otter Classifier with noisy links (N) and no information flow control (C) achieves comparable or even better performance than when we cherry-pick links and control the flow of information (Table 3). Slight variations indicate GNN embeddings’ resilience to irrelevant noisy triples for downstream tasks. Incorporating regression objectives and information flow control doesn’t seem to impact generalization or significantly improve benchmark results.
@@
The base of the Otter-Knowledge is to leverage the pre-trained representations, before enhancing them with additional knowledge. Thus, the initial embeddings computed for the SMILES and sequences have an impact in the results. Table 5 shows the results of Otter-Knowledge using a ClassifierHead, trained on Otter-UBC for 25 epochs with different drug initial representations. We can see that the MolFormer does not give superior results compared to Morgan-fingerprint, there is room for improvement regarding learned representation over simple fingerprint-based approaches for small molecules.
@@
Table 5: Impact of different modalities for drugs on the Otter-UDB datasets
@@ 5 Related work
Representation learning focuses on encoding essential information about entities, such as proteins or molecules, into low-dimensional tensors. Self-supervised algorithms using language models (LMs) have achieved remarkable success in learning protein and molecule representations by training on extensive datasets of protein sequences or linear serialization of small molecules, such as SMILES. State of the art examples of transformer-based protein language models (PLMs) are TAPE [22], ProteinLM [36], ProteinBERT [2], ESM [25], Prottrans [6], and MSA [23]. They are typically trained on masked reconstruction – they learn the likelihood that a particular amino acid appears in a sequence context. Because the probability that a residue will be conserved or not across related sequences is intrinsically tied to its biological role, existent PLMs can capture co-evolutionary and inter-residue contact information [22, 25], and have shown impressive performance on various tasks, such as predicting protein structure [15] and function [22]. Regarding small molecules, their molecular structure can be condensed into linear notations like SMILES or SELFIES. LMs have also been used to interpret these representations, e.g., MolFormer [28], MolBERT [7], SmilesFormer [20] or SELFormer [37]. Both Protein and molecular representations have been fine-tuned using a contrastive learning co-embedding by Conplex [30] achieving good performance in Drug-Target Interaction (DTI) prediction, surpassing state of the art approaches in the TDC-DG leaderboard which evaluates of out-of-domain generalisation [11] and achieving high specificity while detecting false positive bindings in “decoy” datasets like DUDe.
@@
LMs do not consider existent extensive knowledge, in the form of manually curated functional and structural annotations, in human-curated domain datasets and effectively leveraging all this available factual knowledge to enhance representation learning is an open challenge. Nonetheless, prior research indicates that it can improve results in downstream learning tasks. OntoProtein [38] fine-tuned a PLM by reconstructing masked amino acids while minimizing the embedding distance between the contextual representation of proteins and associated gene ontology (GO) functional annotations [3]. For this purpose they built ProteinKG25, a KG consisting of  600k protein sequences and nearly five million triples. Their results show that the representations obtained where useful for classification tasks such as protein-protein interaction type, protein function, and contact prediction; but underperformed in regression tasks like homology, fluorescence, and stability. KeAP [39] claims to explore a more granular token-level approach, where non-masked amino acids iteratively query the associated knowledge tokens to extract helpful information (from GO) for restoring masked amino acids via cross-attention. The training process is guided only by the mask token reconstruction objective, while OntoProtein uses both contrastive learning and masked modelling simultaneously. KeAP, also trained on ProteinKG25 [38] outperforms OntoProtein on 9 dowstream tasks.
@@
Graphs are a natural way to represent molecular interactions, signalling pathways, and disease co-morbidities. They can also be used for representation learning as they allow for the distillation of high-dimensional information about a node’s neighborhood into low-dimensional vector spaces. The training objective of these representations is that similar neighborhoods are embedded close to each other in the vector space. Optimised representations can then be used to train downstream models to predict properties of specific nodes (e.g., protein function), as well as, novel associations between nodes (e.g., drug-target interactions). An overview on graph representation learning in biomedicine can be found in [14]. State-of-the-art approaches have shown that incorporating multiple knowledge sources improves downstream performance. For example, DTiGEMS+ [35] formulates the prediction of DTIs as a link prediction problem in an heterogeneous graph. TxGNN [10] predicts drug indications/contraindications for rare diseases using the PrimeKG heterogeneous and multimodal KG [4]. To the best of our knowledge, we are the first to demonstrate that KGs, derived from various sources combining multiple modalities and specifically tailored to augment protein and SMILES sequences for representation learning, improved DTI downstream tasks, when evaluated using discovery benchmarks [11] wherein a majority of entities in the test data are unseen in the training data.
@@ 6 Conclusion and future work
In this paper we present multi-modal knowledge-enhanced representation learning for proteins and molecules, and we show its relevance and benefits for some drug discovery tasks. Our study lays the foundation for future investigations in this area with the release of multimodal KGs selectively integrating data from several datasets, and pretrained models constructed with our KGs. Our models can be utilized to acquire representations specifically tailored for drug discovery applications. Furthermore, they can serve as benchmarks for comparing and evaluating against other representation learning techniques. We have also publicly released a standard evaluation framework for assessing pretrained representations for prediction of drug-target binding affinity. We extensively analyzed different representation learning methods on three established drug-target binding datasets, and found that our approach outperformed existing methods, achieving state-of-the-art results on the TDC DG dataset.
Numerous unresolved research directions remain unexplored. Firstly, the inclusion of additional modalities, such as the 3D structure of molecules or proteins, can provide valuable insights for representation learning. Secondly, a substantial challenge lies in effectively handling a vast number of datasets, where aligning them into a single multimodal KG is not a straightforward task. Developing a learning approach capable of accommodating the dynamic input schema from diverse sources is a crucial problem to address. Finally, assessing how representations generalize across tasks, and developing robust learning methods for representation generalization across multiple tasks under changing data distributions, remain as key research goals.
@@ 7 Limitations
Due to license restriction of some datasets we only release the multimodal KGs with non-commercial licenses and the pretrained models that were built on these KGs. The KGs do not include 3D structure of proteins/drugs which can be an interesting modalities for future work.
@@ Appendix A Appendix
A.1 Experimental settings
@@
The downstream models comprise of two parallel networks. In the first one, the initial embeddings of drugs and proteins, obtained using the Morgan-fingerprint and ESM-1b models, are concatenated. These embeddings undergo a transformation through two linear layers and RELU activation before predicting the binding affinity. The second architecture involves the concatenation of GNN embeddings of drugs and proteins. Similarly, these embeddings are transformed using two linear layers and RELU activation, followed by predicting the binding affinity.
The results presented in Table 3 rely on pretrained models that have restricted the information they can access for Protein/Drug entities. These entities are only allowed to utilize protein sequences and smiles. Inference tasks follow the pretraining tasks to acquire pretrained GNN embeddings for drugs/proteins in the downstream tasks. Even so Table 4 shows that information flow controls to drugs/proteins is not necessary to achieve good performance on the downstream tasks.
There are two sets of hyper-parameters, one concerning GNN models training and the other concerning donwstream model training. The learning rate of the first set of hyper-parameters was chosen in the range 0.01-0.00001 (log scale) to optimize the loss of the validation split of the pretraining tasks.
The learning rate of the second set of hyper-parameters for the downstream tasks were chosen in the range 0.0005, 0.00005, 0.000005 based on the Pearson correlation score observed on the validation dataset of the TDC-DG dataset. For DAVIS and KIBA where the data is small (has no validation split), we have discussed how to choose the learning rate in Figure 2 and Figure 3. The performance of the baseline and the proposed approach in the paper varied with different values of learning rate, but we always observed better results with the proposed approach in this paper compared to the baseline without knowledge enhanced representation.
@@ A.2 Model release
Table 7 provides a summary of the models. Each model’s name reflects the dataset it was trained on and the scoring function employed during pretraining. For instance, the model otter_ubc_transe denotes the one trained on the Uniprot, BindingDB, and ChemBL (Otter-UBC) dataset using the TransE scoring function. All models underwent training for either 35 epochs or within a maximum time frame of 24 hours.
@@
otter_ubc_transe Otter-UBC TransE MIT
otter_ubc_distmult Otter-UBC DistMult MIT
otter_ubc_classifier Otter-UBC Classifier MIT
otter_dude_transe Otter-DUDe TransE MIT
otter_dude_distmult Otter-DUDe DistMult MIT
otter_dude_classifier Otter-DUDe Classifier MIT
otter_primekg_transe Otter-PrimeKG TransE MIT
otter_primekg_distmult Otter-PrimeKG DistMult MIT
otter_primekg_classifier Otter-PrimeKG Classifier MIT
otter_stitch_transe Otter-Stitch TransE MIT
otter_stitch_distmult Otter-Stitch DistMult MIT
otter_stitch_classifier Otter-Stitch Classifier MIT